OpenStudy (anonymous):

9. Find the solutions of the inequality. Show your work. 4b – 3 > –1

OpenStudy (squirrels):

Okay so since youre new, this is how to you tag people. @MrsPurvis2010

OpenStudy (anonymous):

oh ok thank you

OpenStudy (squirrels):

Anyways how to find the equality. What you wanna do is find 'b'. So lets try to isolate b. \[4b-3>-1\] Now to isolate 'b' lets add 3 to both sides. \[4b>2\]

OpenStudy (squirrels):

3 gets canceled out, leaving us with that equation. Do you know what to do now?

OpenStudy (anonymous):

not really

OpenStudy (squirrels):

But did you understand what I right now?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

but still confused

OpenStudy (squirrels):

Ok so we are trying to find the number for 'b' right?

OpenStudy (squirrels):

Ill show you later, anyways lets continue the equation

OpenStudy (anonymous):

yes its inequality

OpenStudy (squirrels):

\[4b>2\]Now we divide 4 to both sides, to get 'b' by itself. That way we'd get the answer. \[b=\frac{ 1 }{ 2 } or .5\]

OpenStudy (anonymous):

what do i divide 4 too

OpenStudy (anonymous):

i have to show my work

OpenStudy (anonymous):

sorry still confused can you start from the beginning please

OpenStudy (squirrels):

You divide 4 to both sides of the equation. \[\frac{ 4b }{ 4 }>\frac{ 2 }{ 4 }\] \[b=.5\]

OpenStudy (anonymous):

alright i get that but how do i break down 4b-3>-1 I have to show all the work

OpenStudy (squirrels):

you add 3 from both sides of the equation

OpenStudy (squirrels):

\[4b-3+3>-1+3\]

OpenStudy (anonymous):

4b-3>-1 +3 +3 6 2

OpenStudy (squirrels):

No -3+3=0 making that cancel out

OpenStudy (anonymous):

-1+3 +2

OpenStudy (anonymous):

-1+3=2

OpenStudy (squirrels):

Yes

OpenStudy (anonymous):

now what

OpenStudy (squirrels):

thats the answer :P b=2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

if i need more help would you help me

OpenStudy (squirrels):

i will try. :~)

OpenStudy (anonymous):

k